OpenStudy (anonymous):

17. Twelve cans of peaches are placed into each box. Write a function rule for the number of boxes needed for c cans. How many boxes are needed for 1440 cans?

OpenStudy (mathstudent55):

Let b = number of boxes Let c = number of cans b = c/12 For c = 1440 b = 1440/12 b = 120 Answer: 120 boxes

OpenStudy (anonymous):

Thank you for your help

OpenStudy (mathstudent55):

You're welcome

OpenStudy (help!!!!):

120

OpenStudy (anonymous):

Write an equation for the relationship. 6 + 3 * 3